Given that the point (-36,-77) is on the terminal side of an angle, theta, find these functions...sin(theta)? cos(theta)? csc(theta)? and sec(theta)?
Answer by lwsshak3(6498)

Given that the point (-36,-77) is on the terminal side of an angle, theta, find these functions...sin(theta)? cos(theta)? csc(theta)? and sec(theta)?
**
Given info places reference angle in quadrant III. You are working with a right triangle with opposite side=-77, adjacent side=-36 and hypotenuse=85
Trig functions:
sin(theta)=-77/85
cos(theta)=-36/85
csc(theta)=-85/77
sec(theta)=-85/36